OpenStudy (kittiwitti1):

CALCULUS I | Evaluate the following integrals by interpreting in terms of areas (geometric approach)

OpenStudy (kittiwitti1):

\[\int_{-1}^{4}(2x-1)dx\]\[\int_{0}^{4}(2-\sqrt{16-x^2})dx\]

OpenStudy (kainui):

I suggest graphing them if you're unfamiliar with what these shapes are so that you can recognize them in the future.

OpenStudy (kittiwitti1):

#1: finished. I need answer checks. #2: I cannot graph it properly. Should I split up the integral?

OpenStudy (kainui):

Yeah you can split up the integral that's a good idea, then just add up the area of the pieces

OpenStudy (kainui):

I think specifically this might be the hardest part to recognize, but if you know the area of this from before calculus you'll be set.

OpenStudy (kittiwitti1):

My professor said not to, and mentioned something about a quarter circle. I assume that is the second half, but I cannot figure out how to do it without this --\[\int{f(a)}-\int{f(b)}\]

OpenStudy (salty):

For the second one y=2-sqrt{16-x^2} (y-2)^2=16-x^2 Try converting this equation into standard form of circle and find its radius and center Its graph will be a semicircle

OpenStudy (kainui):

What do you mean you were told this? Graph it to see it for yourself.

OpenStudy (kainui):

If you can't graph it, now's the time to learn

OpenStudy (salty):

it will be a quarter circle for x=0 to x=4

OpenStudy (kittiwitti1):

and that splitting up integrals was unnecessary... Oh, I see. Thank you @Salty

OpenStudy (salty):

Yw